OpenStudy (anonymous):

Consider the function f(x) = 12 x^5 + 60 x^4 - 100 x^3 + 1. f(x) has inflection points at (reading from left to right) x = D, E, and F. 1.) What are the 3 inflection points?

OpenStudy (dumbcow):

set f''(x) = 0 f''(x) = 240x^3 + 720x^2 -600x = 0 --> 120x(2x^2+6x -5) = 0 use quadratic formula

OpenStudy (anonymous):

x= 0 x= 1/2(-3(sqrt(19)) = -3.6794 x= 1/2(sqrt(19)-3) = 0.67945 @dumbcow

OpenStudy (dumbcow):

yes